The demand curve for product X is given by QX = 50 − 2PX. How much consumer surplus do consumers receive when PX = $5?

1
Expert's answer
2022-01-23T15:40:11-0500

If PX = $5, then:

Qx = 50 - 2×5 = 40 units.

Consumer surplus is:

CS = 0.5×40×(50 - 5) = $900.